A quality control engineer at a potato chip company tests the bag filling machine by weighing bags of potato chips. Not every bag contains exactly the same weight. But if more than 16% of bags are over-filled then they stop production to fix the machine. They define over-filled to be more than 1 ounce above the weight on the package.

The engineer weighs 105 bags and finds that 26 of them are over-filled. He plans to test the hypotheses H0: p = 0.16 versus Ha: p > 0.16. What is the test statistic?

z =

(Please round your answer to two decimal places if necessary.)
Answer #1

solution

\hat p = x / n = 26/105==0.248

P0 = 0.16

1 - P0 = 1 - 0.84

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.248 - 0.16/ [\sqrt(0.16*084) /105 ]

z= 2.46